Imagine you borrow $800 from your roommate, agreeing to pay her back $800 plus 11 percent nominal interest in one year. Assume inflation over the life of the contract is expected to be 4.08 percent. What is the total dollar amount you will have to pay her back in a year? What approximate percentage of the interest payment is the result of the real rate of interest?
Answer #1

Dollar amount paid back in 1 year = 800+ 800*11% = 888
Real rate of Interest = (1+11%)/(1+4.08%) -1 = 6.6487%
Real part of Interest = 800*(6.6487%) = 53.19
Percentage of real part in interest = 53.19/88 = 60.44%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well